an airplane accelerates down a runway at 3.20m/s2 for 32.8s until id finally lifts off the ground. Determine the distance traveled before the takeoff.
1
Expert's answer
2018-03-13T11:04:07-0400
Normally, airplane accelerates from the rest. Therefore, its initial speed equals to 0. In such conditions, the distance travelled depends only on acceleration and time:

D=(a×t^2)/2=(3.20 m/s^2 ×〖(32.8 s)〗^2)/2=1721.34 m